OpenStudy (anonymous):

What is the angle of rotation for a regular octagon? (Note: Be sure to type out the word degrees, instead of the symbol.)

OpenStudy (anonymous):

The way I always think of this is to imagine walking along the base (ccw) and then I turn left by what angle? Answer: 360/n = 45 because after n turns I must end up facing how I started The other way to think of it is to note that the sum of interior angles is (n-2) 180 = 1080 1080/n = 135 And the external angles are the supplement of this, that is, 45